OpenStudy (anonymous):

To reduce the following rational expression to lowest terms, you will have to use factoring by grouping. Be sure to factor each numerator and denominator completely before dividing out any common factors. (Problem in comments...)

OpenStudy (anonymous):

\[\frac{ xy+8x+5y+40 }{ xy+8x+6x+48 }\]

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

use the given hint, start by factoring numerator and denominator separately

OpenStudy (anonymous):

but it's kind of already there lol 8*5=40 & 8*6=48

OpenStudy (anonymous):

but how would i put it?

OpenStudy (tkhunny):

Every once in a while, factoring by grouping manages to work. (xy+8x)+(5y+40) x(y+8)+5(y+8) (x+5)(y+8) You do the denominator.

OpenStudy (anonymous):

(xy+8x)+(6x+48) x(y+8)+6(y+8) (x+6)(y+8) \[\frac{ (x+5)(y+8) }{ (x+6)(y+8) }\] both y+8s cancel out \[\frac{ (x+5) }{ (x+6) }\]